OpenStudy (anonymous):

The graph of f(x) = (ax+b)/(x-1)(x-4) has a horizontal tangent line at (2,-1). Find a and b.

OpenStudy (anonymous):

f(x) = (ax+b)(x^2 -5x + 4)^-1 f = ax + b f' = a g = (x^2 - 5x + 4)^-1 g' = -1(x^2 - 5x + 4)^-2(2x - 5)

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

I am going to carry on using the product rule..